Hi I need help with this problem: The directions say to find the solution by the addition method.
-x+y=2
x+y=4
If you could show how it is done I would appreciate it. Thank you..

Answer by WannabeCAgirl83(35)

1) - x + y = 2
2) x + y = 4

0 + 2y = 6 → I added equation 1) to equation 2): -x+x=0, y+y=2y, 2+4=6
2y = 6 | :2 on both sides
y = 3 → Now you subsitute y with 3 in equation 2):

x + y = 4
x + 3 = 4 | -3 on both sides
x = 1 → To varify this solution you do the same in equation 1):

-x + y = 2
-x + 3 = 2 | -3 on both sides
-x = -1 | : -1
x = 1 → The solution is varified.

You're welcome. :-)